How much potassium hydroxide is needed to make 1.00 liter of a 0.5 Molar solution of potassium hydroxide ?
1
Expert's answer
2020-05-10T09:20:00-0400

Solution:

C(KOH) = n(KOH)/V(solution)

n(KOH) = C(KOH)*V(KOH) = 0.5 M* 1L =

= 0.5 mol

m(KOH)= n(KOH)*M(KOH) = 0.5 mol*56 g/mol = 28 g